# MapReduce LoRA: Advancing the Pareto Front in Multi-Preference Optimization for Generative Models

### v2

@@ 4.1 Experimental Setup
Models and datasets. For Text-to-Image generation, we use Stable Diffusion 3.5 Medium (SD 3.5 M) [sd35m_hf] and FLUX.1-dev [flux24] as base models. We adopt the same datasets as Flow-GRPO [flow_grpo_arxiv'25]: GenEval [GenEval_NeurIPS'23], PickScore [PickScore_NeurIPS'23], and OCR [paddleocr_arxiv'25]. We compare against CaPO [CaPO_CVPR'25], Flow-GRPO individual experts [flow_grpo_arxiv'25], multi-objective RL with mixed data (MORL-D) and with mixed data plus mixed rewards (MORL-DR), and Rewarded Soup [reward_soup_nips'23]. The merging iteration is 4 for all reported results if not specified. For Text-to-Video generation, we use the HunyuanVideo backbone [hunyuanvideo_arxiv'25]. We follow the Dance-GRPO [dancegrpo_arxiv25] GitHub recipe (optimizer, GRPO grouping, scheduler, data sampling), but apply parameter-efficient LoRA instead of full fine-tuning. We compare to Dance-GRPO individual experts [dancegrpo_arxiv25]. Because the GitHub configuration differs from the paper, we report reproduced baselines trained on the released data for fairness. The merging iteration is 3 for all reported results if not specified. For Language tasks, we use llama2-7B [llama2_arxiv23] as base model, following Bone Soup [bonesoup_acl25] setup. We first train the Llama2-7B [llama2_arxiv23] backbone via supervised fine-tuning (SFT) using the same datasets and hyperparameters as Bone Soup [bonesoup_acl25], then train LoRA experts with PPO [ppo_arxiv17] under the same settings of datasets and training hyperparameters. The merging iteration is 3 for all reported results if not specified. For implementation details, please refer to the Supplement C.
Reward models. For Text-to-Image, we train with three rewards: detection-based text-image alignment (GenEval [GenEval_NeurIPS'23]), human preference (PickScore [PickScore_NeurIPS'23]), and text rendering quality (OCR [paddleocr_arxiv'25]). For evaluation, we additionally report clip-based text-image alignment (VQAScore [VQAScore_ECCV24]), human preference (MPS [MPS_CVPR'24]), and aesthetic quality (VILA [VILA_CVPR'23]). For Text-to-Video, we use two rewards: Visual Quality (VQ) and Motion Quality (MQ) from VideoAlign [videoalign_arxiv25]. For the Language task, we employ four rewards across two tasks: Reddit Summary (Faithful and Preference1) and Helpful Assistant (Helpful and Harmless).
@@ 4.2 Qualitative Results
Fig. 2 shows iterative improvements from MapReduce LoRA and the controllability of RaTE (see also Fig. 5). Compared to the base model and Rewarded Soup [reward_soup_nips'23] (Fig. 4), MapReduce LoRA better handles challenging cases, including uncommon spatial relations (e.g., an elephant below a horse) and high-quality rendering of small text. For more qualitative results on Text-to-Image and Text-to-Video, please refer to the Supplement E.1 and E.2.
@@ 4.3 Quantitative Results
MapReduce LoRA. Table 1 reports Text-to-Image performance on GenEval [GenEval_NeurIPS'23], PickScore [PickScore_NeurIPS'23], OCR [paddleocr_arxiv'25], VQAScore [VQAScore_ECCV24], MPS [MPS_CVPR'24], and VILA [VILA_CVPR'23] using SD 3.5 M [sd35m_hf] and FLUX.1-dev [flux24]. MapReduce LoRA improves both in-domain and out-of-domain metrics. Relative to CaPO [CaPO_CVPR'25], it delivers larger gains across the four CaPO metrics. On CaPO’s in-domain metrics (VQAScore [VQAScore_ECCV24], MPS [MPS_CVPR'24], VILA [VILA_CVPR'23]), which are out-of-domain for us, our improvements are 2.78%, 4.89%, and 3.85%, exceeding CaPO’s 0.66%, 1.42%, and 2.59%. On CaPO’s out-of-domain metric (GenEval [GenEval_NeurIPS'23]), our improvement is 31.88% versus CaPO’s 4.41%. Flow-GRPO individual experts [flow_grpo_arxiv'25], each tuned to a single reward, substantially boost the targeted metric but transfer poorly to others. MORL-D and MORL-DR can be unstable under conflicting objectives; for instance, PickScore often collapses during joint training and is dominated by other rewards (e.g., OCR). In contrast, MapReduce LoRA maintains competitive performance across all targeted rewards.
Table 2 reports Text-to-Video results. We evaluate 1,024 held-out prompts (fixed once). For each prompt, we generate four samples using a fixed set of seeds shared across models and report mean Visual Quality (VQ) and Motion Quality (MQ) from VideoAlign [videoalign_arxiv25]. MapReduce LoRA improves both VQ and MQ (+48.09% and +89.96%) over individually trained experts (+34.55% and +74.10%).
In Figs. 7 and 8, we compare MapReduce LoRA against Llama2 (base) [llama2_arxiv23], Llama2 after SFT, Rewarded Soup [reward_soup_nips'23], and Bone Soup [bonesoup_acl25]. MapReduce LoRA achieves state-of-the-art performance on both tasks, underscoring its cross-modal generality.
@@
HunyuanVideo [hunyuanvideo_arxiv'25] x 3.25 0.95
Individual Experts (DanceGRPO [dancegrpo_arxiv25]) VQ 4.37 (+34.55%) 1.20 (+26.41%)
@@
Rewarded Soup [reward_soup_nips'23] VQ+MQ 4.13 (+27.37%) 1.43 (+50.42%)
@@
Reward-aware token embedding (RaTE). Tables 1 and 3 report RaTE performance on GenEval [GenEval_NeurIPS'23], PickScore [PickScore_NeurIPS'23], and OCR [paddleocr_arxiv'25] using SD 3.5 M [sd35m_hf]. At inference, RaTE appends a trained control token to the end of the prompt based on the specified preference. RaTE enables per-reward control and jointly improves multiple rewards. In Table 1, token gains are orthogonal to transformer tuning: RaTE adds 4.20%, 1.03%, and 2.17% on GenEval [GenEval_NeurIPS'23], PickScore [PickScore_NeurIPS'23], and OCR [paddleocr_arxiv'25], respectively. In Table 3, when using all three tokens, the first appended token accounts for most of the gain. Token control is effective on Stable Diffusion [sd35m_hf, ti_iclr'23], which uses explicit cross-attention for text conditioning. In contrast, FLUX.1-dev [flux24] performs joint text–image sequence modeling, causing a modified token embedding to perturb both text and image tokens across layers, which makes reward-specific token control substantially less stable. We leave this for future exploration.
Table 3: Reward-aware Token Embedding results. <<GE>>, <<PS>>, and <<OCR>> denote tokens trained on GenEval [GenEval_NeurIPS'23], PickScore [PickScore_NeurIPS'23], and OCR [paddleocr_arxiv'25], respectively. Gray indicates not trained on that reward; bold indicates the best among variants.
Method GenEval [GenEval_NeurIPS'23] PickScore [PickScore_NeurIPS'23] OCR [paddleocr_arxiv'25]
SD 3.5 M [sd35m_hf] 0.68 21.784 0.601
@@ 4.4 Ablation Study
[MapReduce LoRA] Effect of merging iterations under a fixed training budget. With the total training steps fixed (Fig. 6), we compare merging iterations $k\in\{1,4,10\}$. Increasing $k$ consistently improves final performance and reduces degradation from merging multiple experts. For $k=10$ relative to $k=4$ and $k=1$, the gains are +1.12% and +11.84% on GenEval [GenEval_NeurIPS'23], +0.57% and +1.56% on PickScore [PickScore_NeurIPS'23], and +0.03% and +10.72% on OCR [paddleocr_arxiv'25].
[RaTE] Effect of appended token count. Table 4 reports RaTE performance as we vary the number of appended tokens. In the table, <<RaTE>> denotes the metric-specific token: <<GE>> for GenEval [GenEval_NeurIPS'23], <<PS>> for PickScore [PickScore_NeurIPS'23], and <<OCR>> for OCR [paddleocr_arxiv'25]. Performance peaks differ by reward: GenEval saturates at 2–3 tokens (0.78), PickScore peaks at 1 (22.052), and OCR peaks at 3 (0.635).
@@
Method GenEval [GenEval_NeurIPS'23] PickScore [PickScore_NeurIPS'23] OCR [paddleocr_arxiv'25]
SD 3.5 M [sd35m_hf] 0.68 21.784 0.601
@@ Appendix A MapReduce LoRA on Language Tasks
Beyond Text-to-Image and Text-to-Video, we evaluate MapReduce LoRA on language tasks to demonstrate cross-modal generality. Following the Bone Soup [bonesoup_acl25] setup, we consider two tasks, each with two rewards: (i) Reddit Summary (Faithful, Preference1) and (ii) Helpful Assistant (Helpful, Harmless). We first train the Llama2-7B [llama2_arxiv23] backbone via supervised fine-tuning (SFT) using the same datasets and hyperparameters as Bone Soup [bonesoup_acl25], then train LoRA experts with PPO [ppo_arxiv17] under the same settings of datasets and training hyperparameters. In Figs. 7 and 8, we compare MapReduce LoRA against Llama2 (base) [llama2_arxiv23], Llama2 after SFT, Rewarded Soup [reward_soup_nips'23], and Bone Soup [bonesoup_acl25]. MapReduce LoRA achieves state-of-the-art performance on both tasks, underscoring its cross-modal generality.
Figure 7: Language Task: Reddit Summary. Left: MapReduce LoRA outperforms Rewarded Soup [reward_soup_nips'23] and Bone Soup [bonesoup_acl25] (reproduced from Fig. 5(c) in [bonesoup_acl25]) on both rewards. Right: MapReduce LoRA improves progressively across merge iterations.
Figure 8: Language Task: Helful Assistant. Left: MapReduce LoRA outperforms Rewarded Soup [reward_soup_nips'23] and Bone Soup [bonesoup_acl25] (reproduced from Fig. 5(a) in [bonesoup_acl25]) on both rewards. Right: MapReduce LoRA improves progressively across merging iterations.
@@ C.1 Training Configuration
For each reward-aware token embedding, it contains 3 embeddings corresponding to 3 text encoders within SD 3.5 M [sd35m_hf]. The embedding is trained on the GenEval, PickScore, and OCR datasets, individually. The teacher model is trained with LoRA via GRPO [grpo_arxiv'24] with 4,200 steps on GenEval [GenEval_NeurIPS'23], 4,500 steps on PickScore [PickScore_NeurIPS'23] and 1,600 steps on OCR [paddleocr_arxiv'25]. The training batchsize for GenEval is 512, PickScore is 768 and OCR is 768. We use the AdamW optimizer [adamw_iclr'19] with learning rate 3e-4, $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999, weight decay = 1e-4 and no warmup schedule; no KL penalty is applied.
@@
  • Text-to-Image: We train all models using AdamW optimizer [adamw_iclr'19] with learning rate 3e-4, $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999, weight decay = 1e-4, and no warmup schedule. All experiments use 32 GPUs with global batch size 576.
@@
      SD 3.5 M [sd35m_hf]: We follow the training configuration in Flow-GRPO [flow_grpo_arxiv'25], where only the KL ratio $\beta$ is different. $\beta_{\text{GenEval}}$ = 0.04, $\beta_{\text{PickScore}}$ = 0.01, and $\beta_{\text{OCR}}$ = 0.04. The sampling timestep $T$ is 10 and the evaluation timestep $T$ = 40. The GRPO group size $G$ is 24, the resolution is 512, and the LoRA settings are $\alpha$ = 64 and $r$ = 32. LoRA is applied to all attention layers (attn.add_q_proj, attn.add_k_proj, attn.add_v_proj, attn.to_add_out, attn.to_q, attn.to_k, attn.to_v, attn.to_out.0). The per-GPU batch sizes are 6, 9, and 9 with gradient accumulation steps of 3, 2, and 2 for GenEval, PickScore, and OCR, respectively. We train GenEval, PickScore and OCR for {4,100, 1,200, 1,000, 1,100}, {4,500, 1,400, 1,000, 1,000}, and {1,600, 1,300, 900, 1,100} steps, respectively; the bracketed numbers denote merge iterations 0-3.
@@
      FLUX.1-dev [flux24]: The only task-dependent change is the KL ratio $\beta$ ($\beta_{\text{GenEval}}{=}0.04$, $\beta_{\text{PickScore}}{=}0$, $\beta_{\text{OCR}}{=}0.04$). The sampling timestep $T$ is 6 and the evaluation timestep $T$ = 28. The GRPO group size $G$ is 24, the resolution is 512, and the LoRA settings are $\alpha$ = 128 and $r$ = 64. LoRA is applied to attention layers and feed-forward network layers (attn.to_q, attn.to_k, attn.to_v, attn.to_out.0, attn.add_q_proj, attn.add_k_proj, attn.add_v_proj, attn.to_add_out, ff.net.0.proj, ff.net.2, ff_context.net.0.proj,
@@
  • Text-to-Video: We train the HunyuanVideo [hunyuanvideo_arxiv'25] with the same configuration as the DanceGRPO [dancegrpo_arxiv25] but apply LoRA instead of full finetuning. The LoRA settings are $\alpha$ = 64 and $r$ = 32. LoRA is applied to attention layers “attn.to_q, attn.to_k, attn.to_v” in single_transformer_blocks (40 blocks), “attn.to_q, attn.to_k, attn.to_v, attn.to_out.0, attn.to_add_out, attn.add_q_proj, attn.add_k_proj, attn.add_v_proj” in transformer_blocks (20 blocks). We train VQ and MQ for {200, 140, 100}, {195, 145, 115} steps, respectively; the bracketed numbers denote merge iterations 0-2.
@@
The hyperparameters for MORL are identical to the MapReduce LoRA mentioned above. During each epoch, batches are sampled from GenEval [GenEval_NeurIPS'23], PickScore [PickScore_NeurIPS'23], and OCR [paddleocr_arxiv'25] datasets according to specified ratios (default 1:1:1), where each individual batch contains samples from a single source. During evaluation, each data source is independently evaluated using its corresponding reward model to provide per-task performance metrics. The training steps are comparable to the overall training steps of MapReduce LoRA for fairness. There are two variations of MORL in our experiments, and the key difference lies in the data mixture and reward mixture strategy:
@@
  • Data mixture and reward mixture (MORL-DR): Samples from GenEval [GenEval_NeurIPS'23] and OCR [paddleocr_arxiv'25] are scored by both their task-specific reward and PickScore [PickScore_NeurIPS'23]; we take the average as the final reward (e.g., for GenEval, $(r_{\text{geneval}}+r_{\text{pickscore}})/2$). In contrast, samples from the PickScore dataset are evaluated only with PickScore. This choice reflects dataset constraints: GenEval requires a structured prompt format, whereas OCR requires prompts that contain text for a valid evaluation. PickScore serves as a shared aesthetic quality anchor across tasks. Overall, this protocol better aligns with the multi-objective setting by applying multiple rewards in a dataset-agnostic manner (where applicable), rather than restricting evaluation to dataset-specific rewards.

### v3

@@ 4.1 Experimental Setup
Models and datasets. For Text-to-Image generation, we use Stable Diffusion 3.5 Medium (SD 3.5 M) [3] and FLUX.1-dev [17] as base models. We adopt the same datasets as Flow-GRPO [23]: GenEval [11], PickScore [15], and OCR [6]. We compare against CaPO [18], Flow-GRPO individual experts [23], multi-objective RL with mixed data (MORL-D) and with mixed data plus mixed rewards (MORL-DR), and Rewarded Soup [33]. The merging iteration is 4 for all reported results if not specified. For Text-to-Video generation, we use the HunyuanVideo backbone [16]. We follow the Dance-GRPO [41] GitHub recipe (optimizer, GRPO grouping, scheduler, data sampling), but apply parameter-efficient LoRA instead of full fine-tuning. We compare to Dance-GRPO individual experts [41]. Because the GitHub configuration differs from the paper, we report reproduced baselines trained on the released data for fairness. The merging iteration is 3 for all reported results if not specified. For Language tasks, we use llama2-7B [38] as base model, following Bone Soup [40] setup. We first train the Llama2-7B [38] backbone via supervised fine-tuning (SFT) using the same datasets and hyperparameters as Bone Soup [40], then train LoRA experts with PPO [35] under the same settings of datasets and training hyperparameters. The merging iteration is 3 for all reported results if not specified. For implementation details, please refer to the Supplement C.
Reward models. For Text-to-Image, we train with three rewards: detection-based text-image alignment (GenEval [11]), human preference (PickScore [15]), and text rendering quality (OCR [6]). For evaluation, we additionally report clip-based text-image alignment (VQAScore [21]), human preference (MPS [43]), and aesthetic quality (VILA [14]). For Text-to-Video, we use two rewards: Visual Quality (VQ) and Motion Quality (MQ) from VideoAlign [24]. For the Language task, we employ four rewards across two tasks: Reddit Summary (Faithful and Preference1) and Helpful Assistant (Helpful and Harmless).
@@ 4.2 Qualitative Results
Fig. 2 shows iterative improvements from MapReduce LoRA and the controllability of RaTE (see also Fig. 5). Compared to the base model and Rewarded Soup [33] (Fig. 4), MapReduce LoRA better handles challenging cases, including uncommon spatial relations (e.g., an elephant below a horse) and high-quality rendering of small text. For more qualitative results on Text-to-Image and Text-to-Video, please refer to the Supplement E.1 and E.2.
@@ 4.3 Quantitative Results
MapReduce LoRA. Table 1 reports Text-to-Image performance on GenEval [11], PickScore [15], OCR [6], VQAScore [21], MPS [43], and VILA [14] using SD 3.5 M [3] and FLUX.1-dev [17]. MapReduce LoRA improves both in-domain and out-of-domain metrics. Relative to CaPO [18], it delivers larger gains across the four CaPO metrics. On CaPO’s in-domain metrics (VQAScore [21], MPS [43], VILA [14]), which are out-of-domain for us, our improvements are 2.78%, 4.89%, and 3.85%, exceeding CaPO’s 0.66%, 1.42%, and 2.59%. On CaPO’s out-of-domain metric (GenEval [11]), our improvement is 31.88% versus CaPO’s 4.41%. Flow-GRPO individual experts [23], each tuned to a single reward, substantially boost the targeted metric but transfer poorly to others. MORL-D and MORL-DR can be unstable under conflicting objectives; for instance, PickScore often collapses during joint training and is dominated by other rewards (e.g., OCR). In contrast, MapReduce LoRA maintains competitive performance across all targeted rewards.
Table 2 reports Text-to-Video results. We evaluate 1,024 held-out prompts (fixed once). For each prompt, we generate four samples using a fixed set of seeds shared across models and report mean Visual Quality (VQ) and Motion Quality (MQ) from VideoAlign [24]. MapReduce LoRA improves both VQ and MQ (+48.09% and +89.96%) over individually trained experts (+34.55% and +74.10%).
In Figs. 7 and 8, we compare MapReduce LoRA against Llama2 (base) [38], Llama2 after SFT, Rewarded Soup [33], and Bone Soup [40]. MapReduce LoRA achieves state-of-the-art performance on both tasks, underscoring its cross-modal generality.
@@
HunyuanVideo [16] x 3.25 0.95
Individual Experts (DanceGRPO [41]) VQ 4.37 (+34.55%) 1.20 (+26.41%)
@@
Rewarded Soup [33] VQ+MQ 4.13 (+27.37%) 1.43 (+50.42%)
@@
Reward-aware token embedding (RaTE). Tables 1 and 3 report RaTE performance on GenEval [11], PickScore [15], and OCR [6] using SD 3.5 M [3]. At inference, RaTE appends a trained control token to the end of the prompt based on the specified preference. RaTE enables per-reward control and jointly improves multiple rewards. In Table 1, token gains are orthogonal to transformer tuning: RaTE adds 4.20%, 1.03%, and 2.17% on GenEval [11], PickScore [15], and OCR [6], respectively. In Table 3, when using all three tokens, the first appended token accounts for most of the gain. Token control is effective on Stable Diffusion [3, 10], which uses explicit cross-attention for text conditioning. In contrast, FLUX.1-dev [17] performs joint text–image sequence modeling, causing a modified token embedding to perturb both text and image tokens across layers, which makes reward-specific token control substantially less stable. We leave this for future exploration.
Table 3: Reward-aware Token Embedding results. <<GE>>, <<PS>>, and <<OCR>> denote tokens trained on GenEval [11], PickScore [15], and OCR [6], respectively. Gray indicates not trained on that reward; bold indicates the best among variants.
Method GenEval [11] PickScore [15] OCR [6]
SD 3.5 M [3] 0.68 21.784 0.601
@@ 4.4 Ablation Study
[MapReduce LoRA] Effect of merging iterations under a fixed training budget. With the total training steps fixed (Fig. 6), we compare merging iterations $k\in\{1,4,10\}$. Increasing $k$ consistently improves final performance and reduces degradation from merging multiple experts. For $k=10$ relative to $k=4$ and $k=1$, the gains are +1.12% and +11.84% on GenEval [11], +0.57% and +1.56% on PickScore [15], and +0.03% and +10.72% on OCR [6].
[RaTE] Effect of appended token count. Table 4 reports RaTE performance as we vary the number of appended tokens. In the table, <<RaTE>> denotes the metric-specific token: <<GE>> for GenEval [11], <<PS>> for PickScore [15], and <<OCR>> for OCR [6]. Performance peaks differ by reward: GenEval saturates at 2–3 tokens (0.78), PickScore peaks at 1 (22.052), and OCR peaks at 3 (0.635).
@@
Method GenEval [11] PickScore [15] OCR [6]
SD 3.5 M [3] 0.68 21.784 0.601
@@ Appendix A MapReduce LoRA on Language Tasks
Beyond Text-to-Image and Text-to-Video, we evaluate MapReduce LoRA on language tasks to demonstrate cross-modal generality. Following the Bone Soup [40] setup, we consider two tasks, each with two rewards: (i) Reddit Summary (Faithful, Preference1) and (ii) Helpful Assistant (Helpful, Harmless). We first train the Llama2-7B [38] backbone via supervised fine-tuning (SFT) using the same datasets and hyperparameters as Bone Soup [40], then train LoRA experts with PPO [35] under the same settings of datasets and training hyperparameters. In Figs. 7 and 8, we compare MapReduce LoRA against Llama2 (base) [38], Llama2 after SFT, Rewarded Soup [33], and Bone Soup [40]. MapReduce LoRA achieves state-of-the-art performance on both tasks, underscoring its cross-modal generality.
Figure 7: Language Task: Reddit Summary. Left: MapReduce LoRA outperforms Rewarded Soup [33] and Bone Soup [40] (reproduced from Fig. 5(c) in [40]) on both rewards. Right: MapReduce LoRA improves progressively across merge iterations.
Figure 8: Language Task: Helful Assistant. Left: MapReduce LoRA outperforms Rewarded Soup [33] and Bone Soup [40] (reproduced from Fig. 5(a) in [40]) on both rewards. Right: MapReduce LoRA improves progressively across merging iterations.
@@ C.1 Training Configuration
For each reward-aware token embedding, it contains 3 embeddings corresponding to 3 text encoders within SD 3.5 M [3]. The embedding is trained on the GenEval, PickScore, and OCR datasets, individually. The teacher model is trained with LoRA via GRPO [36] with 4,200 steps on GenEval [11], 4,500 steps on PickScore [15] and 1,600 steps on OCR [6]. The training batchsize for GenEval is 512, PickScore is 768 and OCR is 768. We use the AdamW optimizer [26] with learning rate 3e-4, $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999, weight decay = 1e-4 and no warmup schedule; no KL penalty is applied.
@@
  • Text-to-Image: We train all models using AdamW optimizer [26] with learning rate 3e-4, $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999, weight decay = 1e-4, and no warmup schedule. All experiments use 32 GPUs with global batch size 576.
@@
      SD 3.5 M [3]: We follow the training configuration in Flow-GRPO [23], where only the KL ratio $\beta$ is different. $\beta_{\text{GenEval}}$ = 0.04, $\beta_{\text{PickScore}}$ = 0.01, and $\beta_{\text{OCR}}$ = 0.04. The sampling timestep $T$ is 10 and the evaluation timestep $T$ = 40. The GRPO group size $G$ is 24, the resolution is 512, and the LoRA settings are $\alpha$ = 64 and $r$ = 32. LoRA is applied to all attention layers (attn.add_q_proj, attn.add_k_proj, attn.add_v_proj, attn.to_add_out, attn.to_q, attn.to_k, attn.to_v, attn.to_out.0). The per-GPU batch sizes are 6, 9, and 9 with gradient accumulation steps of 3, 2, and 2 for GenEval, PickScore, and OCR, respectively. We train GenEval, PickScore and OCR for {4,100, 1,200, 1,000, 1,100}, {4,500, 1,400, 1,000, 1,000}, and {1,600, 1,300, 900, 1,100} steps, respectively; the bracketed numbers denote merge iterations 0-3.
@@
      FLUX.1-dev [17]: The only task-dependent change is the KL ratio $\beta$ ($\beta_{\text{GenEval}}{=}0.04$, $\beta_{\text{PickScore}}{=}0$, $\beta_{\text{OCR}}{=}0.04$). The sampling timestep $T$ is 6 and the evaluation timestep $T$ = 28. The GRPO group size $G$ is 24, the resolution is 512, and the LoRA settings are $\alpha$ = 128 and $r$ = 64. LoRA is applied to attention layers and feed-forward network layers (attn.to_q, attn.to_k, attn.to_v, attn.to_out.0, attn.add_q_proj, attn.add_k_proj, attn.add_v_proj, attn.to_add_out, ff.net.0.proj, ff.net.2, ff_context.net.0.proj,
@@
  • Text-to-Video: We train the HunyuanVideo [16] with the same configuration as the DanceGRPO [41] but apply LoRA instead of full finetuning. The LoRA settings are $\alpha$ = 64 and $r$ = 32. LoRA is applied to attention layers “attn.to_q, attn.to_k, attn.to_v” in single_transformer_blocks (40 blocks), “attn.to_q, attn.to_k, attn.to_v, attn.to_out.0, attn.to_add_out, attn.add_q_proj, attn.add_k_proj, attn.add_v_proj” in transformer_blocks (20 blocks). We train VQ and MQ for {200, 140, 100}, {195, 145, 115} steps, respectively; the bracketed numbers denote merge iterations 0-2.
@@
The hyperparameters for MORL are identical to the MapReduce LoRA mentioned above. During each epoch, batches are sampled from GenEval [11], PickScore [15], and OCR [6] datasets according to specified ratios (default 1:1:1), where each individual batch contains samples from a single source. During evaluation, each data source is independently evaluated using its corresponding reward model to provide per-task performance metrics. The training steps are comparable to the overall training steps of MapReduce LoRA for fairness. There are two variations of MORL in our experiments, and the key difference lies in the data mixture and reward mixture strategy:
@@
  • Data mixture and reward mixture (MORL-DR): Samples from GenEval [11] and OCR [6] are scored by both their task-specific reward and PickScore [15]; we take the average as the final reward (e.g., for GenEval, $(r_{\text{geneval}}+r_{\text{pickscore}})/2$). In contrast, samples from the PickScore dataset are evaluated only with PickScore. This choice reflects dataset constraints: GenEval requires a structured prompt format, whereas OCR requires prompts that contain text for a valid evaluation. PickScore serves as a shared aesthetic quality anchor across tasks. Overall, this protocol better aligns with the multi-objective setting by applying multiple rewards in a dataset-agnostic manner (where applicable), rather than restricting evaluation to dataset-specific rewards.